OpenStudy (anonymous):

A population of flies grows according to the function p(x) = 3(2)x, where x is measured in weeks. A local spider has set up shop and consumes flies according to the function s(x) = 2x + 4. What is the population of flies after three weeks with the introduced spide

OpenStudy (anonymous):

Think of how many flies can be produced after 3 weeks (x=3) >> p(x=3) = 6*(3) = 18 with introduced spider that can consume flies as a function of s(x=3) = 2*(3)+4 Therefore, p(x) - s(x) = 18-10 = 8